OpenStudy (anonymous):

Assume that sin x =.8 and sin y =(square root(.75)) X and Y lie between 0 and ((pi)/2) Evaluate sin (x+y) Show all work Please?

OpenStudy (anonymous):

sin(x+y)=sin(x)cos(y)+sin(y)cos(x) sin(x)=.8 therefore cos(x)= \[\sqrt{1-.8^{2}}\] = \[\sqrt{1-.64}=\sqrt{.36}=.6\] sin(y)=root(.75) so \[\cos(y)=\sqrt{1-\sqrt{.75}}=\sqrt{1-.75}=\sqrt{.25}=.5\] both sine and cosine are positive because in quadrant 1. hence \[\sin(x+y)=\sin(x)\cos(y) + \cos(x)\sin(y)=.8\times.5+.6\times \sqrt{.75}\]

OpenStudy (anonymous):

oops typo. \[\cos(y)=\sqrt{1-\sqrt{.75}^2}=\sqrt{1-.75}=\sqrt{.25}=.5\]

OpenStudy (anonymous):

nostalgic?

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

the good old days. before drawing before "legend" before wiggling icons

hero (hero):

Ha! Not anymore, lol